OpenStudy (anonymous):

Of the 600 students in a school, 1/3 are in the band, 1/5 sing in the choir, and 1/4 of the remainder are in the drama club. No student is in more than one of these activities. How many of the school's students do not participate in any of these activities?

OpenStudy (anonymous):

130

OpenStudy (anonymous):

if total student =t=600 600/3=200 are in band 600/5=120 sing choir remainder =600-200-120=280 so 280/4=70 are in drama club total participents will be then 200+120+70=390 =>600-390=210 are non pariticipents abtucally